# Serving Large Language Models on Huawei CloudMatrix384

### v2

@@ 5.1. Experimental Setup
Our evaluation is conducted on a Huawei CloudMatrix384 supernode, provided by the ModelArts Lite (Cluster Mode) service in Huawei Cloud. Specifically, we utilize a configuration comprising 256 Ascend 910C NPUs and their associated host Kunpeng CPUs from a single CloudMatrix384. The serving system consists of an LLM inference engine optimized by Huawei and SiliconFlow [^2] https://www.siliconflow.com/ together, deployed with the requisite Huawei CANN software packages. The elastic memory service (EMS) in Huawei Cloud, providing distributed caching capabilities as detailed in §4.4, is pre-deployed across the allocated compute nodes. The entire deployment adheres to our proposed peer-to-peer serving architecture with PDC disaggregation (§4.1), with the following specific configurations for each logical cluster:
Decode Cluster: We deploy a single decode instance utilizing 160 Ascend 910C NPUs (across 20 compute nodes, yielding 320 NPU dies). This instance employs an expert parallelism degree of 320 (EP320) for the sparse MoE layers. For other components like MLA and dense FFN layers, a data parallelism degree of 320 (DP320) is used across the NPU dies. Within these 320 EP ranks, we deploy one expert instance per rank. The expert configuration comprises 32 copies of the shared expert, 256 distinct router experts, and an additional 32 redundant router experts to facilitate expert parallelism load balancing (EPLB).
Prefill Cluster: The prefill cluster consists of 6 prefill instances, each allocated 16 Ascend 910C NPUs (two compute nodes per instance, yielding 32 NPU dies). In total, the prefill cluster uses 96 NPUs. Each prefill instance employs an expert parallelism degree of 32 (EP32) for sparse MoE layers. MLA components within prefill instances utilize a hybrid parallelism strategy detailed in §4.3.1. For expert deployment within each 32-rank prefill instance, we configure 10 experts per rank, consisting of 1 shared expert, 8 router-selected experts, and 1 redundant router expert for effective EPLB.
@@
This experimental configuration serves as the basis for the accuracy and performance evaluation in subsequent sections. The DeepSeek-R1 model evaluated is the 671B parameter version, which has been quantized to INT8 (§4.5) for execution on the Ascend 910C NPUs.
@@ 5.3. Accuracy
Table 6. Accuracy comparison of DeepSeek-R1 with INT8 quantization on Ascend 910C, the official DeepSeek-R1 API (DeepSeek, 2025), and results reported in the DeepSeek-R1 technical report (DeepSeek-AI et al., 2025) across multiple benchmarks (Results from benchmarks with testing configurations deemed inconsistent have been excluded.).
@@
The comparative accuracy results are presented in Table 6. Overall, our DeepSeek-R1 (INT8) implementation on Ascend 910C demonstrates performance largely comparable to both the official DeepSeek-R1 API and the metrics reported in the original technical paper. This indicates that the INT8 quantization applied for deployment on Ascend 910C effectively preserves the model’s capabilities across a diverse range of tasks.
@@ 5.4.1. Microbatch-based Pipeline
(a) Prefill throughput.
(b) Prefill latency breakdown (4K prompt length).
@@ 5.5. Performance of Operators
Understanding the performance of fundamental computation and communication operators is key to diagnosing system bottlenecks and guiding software optimization efforts. In this subsection, we present a micro-benchmark analysis of critical operators relevant to LLM serving, specifically MoE communication primitives, MLA computations, and general matrix multiplication (GEMM) kernels. We evaluate their performance on the CloudMatrix384 (per Ascend 910C die) and compare them against representative performance on NVIDIA H800 GPUs.
@@ 5.5.2. MLA Operator
Table 8. Utilization of MLA operators on NVIDIA H800 and an Ascend 910C die (CloudMatrix384) in compute-intensive settings (BF16/FP16 precision).
@@
CANN MLA on Ascend 910C die BF16/FP16 65.4
Table 9. Memory bandwidth utilization of MLA operators on NVIDIA H800 and an Ascend 910C die (CloudMatrix384) in memory-intensive settings.
Operator Implementation Achieved Bandwidth (GB/s) Hardware Bandwidth (GB/s, Peak) Utilization
DeepSeek FlashMLA on H800 3,000 3,350 89.6%
CANN MLA on Ascend 910C die 1,346 1,600 84.1%
@@
In memory-intensive settings, the efficiency of utilizing available memory bandwidth is paramount. Table 9 shows this comparison. The DeepSeek FlashMLA on H800 achieves an impressive 3,000 GB/s of memory bandwidth, representing 89.6% utilization of its 3,350 GB/s hardware memory bandwidth. Our CANN MLA implementation on CloudMatrix384 achieves 1,346 GB/s against a hardware memory bandwidth of 1,600 GB/s for the NPU configuration used, resulting in a similarly high utilization of 84.1%.
@@ 5.5.3. GEMM Operator
General Matrix Multiplication (GEMM) is a fundamental compute kernel in virtually all deep learning models. The efficiency of GEMM operations, particularly at lower precisions like INT8, is critical for achieving high inference throughput. We benchmark the performance of INT8 GEMM kernels provided by CANN on a single Ascend 910C die (within the CloudMatrix384 system) across a range of matrix dimensions. The results, detailed in Table 10, showcase achieved compute utilization and the sustained memory bandwidth during these operations. These tests are conducted using common GEMM tiling dimensions (BM $\times$ BN = 128 $\times$ 152), with the operations involving INT8 inputs and BF16 outputs.
Table 10. INT8 GEMM achieved memory bandwidth and compute utilization on an Ascend 910C die (CloudMatrix384) across different configurations, using INT8 inputs and BF16 outputs. Tiling: BM $\times$ BN = 128 $\times$ 152.
@@
As indicated in Table 10, the CANN INT8 GEMM kernels on the Ascend 910C die demonstrate consistently high compute utilization, ranging from 77.4% to 82.7% across various matrix shapes (M, N, K) and group counts. For example, with 4 groups and dimensions M=7168, N=4096, K=8192, the kernel achieves an 82.7% utilization. This high efficiency is maintained across different configurations, indicating robust performance of the INT8 compute units on the Ascend 910C die.
The table also reports the sustained memory bandwidth achieved during these GEMM operations, which ranges from 195 GB/s to 327 GB/s. These values are substantially below the Ascend 910C die’s peak memory bandwidth (1,600 GB/s, as noted in the MLA operator analysis, §5.5.2). This observation, when coupled with the high compute utilization figures, strongly suggests that these INT8 GEMM operations are predominantly compute-bound rather than memory-bandwidth-bound for the tested matrix dimensions. Such a characteristic indicates efficient data reuse within the NPU’s internal cache hierarchy and registers, allowing the compute units to operate at a high fraction of their peak capability without being consistently starved for data transfers from memory.
@@ 7. Conclusion
In this paper, we introduce Huawei CloudMatrix, a next-generation AI datacenter architecture that embodies Huawei’s vision for advanced AI infrastructure. We specifically highlight Huawei CloudMatrix384, the first production-grade implementation of this innovative architectural concept. CloudMatrix384 is an AI supernode engineered to efficiently support large-scale AI workloads, featuring a fully peer-to-peer interconnected hardware design. It integrates 384 Ascend 910C NPUs and 192 Kunpeng CPUs interconnected via an ultra-high-bandwidth, low-latency Unified Bus (UB) network. This unique architecture facilitates dynamic resource pooling, streamlined memory management, and exceptional inter-node communication, effectively addressing the scalability and efficiency challenges common in traditional datacenter architecture.

### v3

@@ 5.1. Experimental Setup
Our evaluation is conducted on a Huawei CloudMatrix384 supernode, provided by the ModelArts Lite (Cluster Mode) service in Huawei Cloud. Specifically, we utilize a configuration comprising 256 Ascend 910 NPUs and their associated host Kunpeng CPUs from a single CloudMatrix384. The serving system consists of an LLM inference engine optimized by Huawei and SiliconFlow [^2] https://www.siliconflow.com/ together, deployed with the requisite Huawei CANN software packages. The elastic memory service (EMS) in Huawei Cloud, providing distributed caching capabilities as detailed in §4.4, is pre-deployed across the allocated compute nodes. The entire deployment adheres to our proposed peer-to-peer serving architecture with PDC disaggregation (§4.1), with the following specific configurations for each logical cluster:
Decode Cluster: We deploy a single decode instance utilizing 160 Ascend 910 NPUs (across 20 compute nodes, yielding 320 NPU dies). This instance employs an expert parallelism degree of 320 (EP320) for the sparse MoE layers. For other components like MLA and dense FFN layers, a data parallelism degree of 320 (DP320) is used across the NPU dies. Within these 320 EP ranks, we deploy one expert instance per rank. The expert configuration comprises 32 copies of the shared expert, 256 distinct router experts, and an additional 32 redundant router experts to facilitate expert parallelism load balancing (EPLB).
Prefill Cluster: The prefill cluster consists of 6 prefill instances, each allocated 16 Ascend 910 NPUs (two compute nodes per instance, yielding 32 NPU dies). In total, the prefill cluster uses 96 NPUs. Each prefill instance employs an expert parallelism degree of 32 (EP32) for sparse MoE layers. MLA components within prefill instances utilize a hybrid parallelism strategy detailed in §4.3.1. For expert deployment within each 32-rank prefill instance, we configure 10 experts per rank, consisting of 1 shared expert, 8 router-selected experts, and 1 redundant router expert for effective EPLB.
@@
This experimental configuration serves as the basis for the accuracy and performance evaluation in subsequent sections. The DeepSeek-R1 model evaluated is the 671B parameter version, which has been quantized to INT8 (§4.5) for execution on the Ascend 910 NPUs.
@@ 5.3. Accuracy
Table 5. Accuracy comparison of DeepSeek-R1 with INT8 quantization on Ascend 910, the official DeepSeek-R1 API (DeepSeek, 2025), and results reported in the DeepSeek-R1 technical report (DeepSeek-AI et al., 2025) across multiple benchmarks (Results from benchmarks with testing configurations deemed inconsistent have been excluded.).
@@
The comparative accuracy results are presented in Table 5. Overall, our DeepSeek-R1 (INT8) implementation on Ascend 910 demonstrates performance largely comparable to both the official DeepSeek-R1 API and the metrics reported in the original technical paper. This indicates that the INT8 quantization applied for deployment on Ascend 910 effectively preserves the model’s capabilities across a diverse range of tasks.
@@ 5.4.1. Microbatch-based Pipeline
(a) Prefill throughput.
(b) Prefill latency breakdown (4K prompt length).
@@ 5.5. Performance of Operators
Understanding the performance of fundamental computation and communication operators is key to diagnosing system bottlenecks and guiding software optimization efforts. In this subsection, we present a micro-benchmark analysis of critical operators relevant to LLM serving, specifically MoE communication primitives, MLA computations, and general matrix multiplication (GEMM) kernels. We evaluate their performance on the CloudMatrix384 (per Ascend 910 die) and compare them against representative performance on NVIDIA H800 GPUs.
@@ 5.5.2. MLA Operator
Table 7. Utilization of MLA operators on NVIDIA H800 and an Ascend 910 die (CloudMatrix384) in compute-intensive settings (BF16/FP16 precision).
@@
CANN MLA on Ascend 910 die BF16/FP16 65.4
Table 8. Memory bandwidth utilization of MLA operators on NVIDIA H800 and an Ascend 910 die (CloudMatrix384) in memory-intensive settings.
Operator Implementation Utilization
DeepSeek FlashMLA on H800 89.6%
CANN MLA on Ascend 910 die 84.1%
@@
In memory-intensive settings, the efficiency of utilizing available memory bandwidth is paramount. Table 8 shows this comparison. The DeepSeek FlashMLA on H800 achieves an 89.6% utilization of its hardware memory bandwidth. Our CANN MLA implementation on CloudMatrix384 achieves a similarly high utilization of 84.1%.
@@ 5.5.3. GEMM Operator
General Matrix Multiplication (GEMM) is a fundamental compute kernel in virtually all deep learning models. The efficiency of GEMM operations, particularly at lower precisions like INT8, is critical for achieving high inference throughput. We benchmark the performance of INT8 GEMM kernels provided by CANN on a single Ascend 910 die (within the CloudMatrix384 system) across a range of matrix dimensions. The results, detailed in Table 9, showcase achieved compute utilization and the sustained memory bandwidth during these operations. These tests are conducted using common GEMM tiling dimensions (BM $\times$ BN = 128 $\times$ 152), with the operations involving INT8 inputs and BF16 outputs.
As indicated in Table 9, the CANN INT8 GEMM kernels on the Ascend 910 die demonstrate consistently high compute utilization, ranging from 77.4% to 82.7% across various matrix shapes (M, N, K) and group counts. For example, with 4 groups and dimensions M=7168, N=4096, K=8192, the kernel achieves an 82.7% utilization. This high efficiency is maintained across different configurations, indicating robust performance of the INT8 compute units on the Ascend 910 die.
The table also reports the sustained memory bandwidth achieved during these GEMM operations, which ranges from 195 GB/s to 327 GB/s. These values are substantially below the Ascend 910 die’s peak memory bandwidth. This observation, when coupled with the high compute utilization figures, strongly suggests that these INT8 GEMM operations are predominantly compute-bound rather than memory-bandwidth-bound for the tested matrix dimensions. Such a characteristic indicates efficient data reuse within the NPU’s internal cache hierarchy and registers, allowing the compute units to operate at a high fraction of their peak capability without being consistently starved for data transfers from memory.
Table 9. INT8 GEMM achieved memory bandwidth and compute utilization on an Ascend 910 die (CloudMatrix384) across different configurations, using INT8 inputs and BF16 outputs. Tiling: BM $\times$ BN = 128 $\times$ 152.
@@ 7. Conclusion
In this paper, we introduce Huawei CloudMatrix, a next-generation AI datacenter architecture that embodies Huawei’s vision for advanced AI infrastructure. We specifically highlight Huawei CloudMatrix384, the first production-grade implementation of this innovative architectural concept. CloudMatrix384 is an AI supernode engineered to efficiently support large-scale AI workloads, featuring a fully peer-to-peer interconnected hardware design. It integrates 384 Ascend 910 NPUs and 192 Kunpeng CPUs interconnected via an ultra-high-bandwidth, low-latency Unified Bus (UB) network. This unique architecture facilitates dynamic resource pooling, streamlined memory management, and exceptional inter-node communication, effectively addressing the scalability and efficiency challenges common in traditional datacenter architecture.